A company has designed a coffee vending machine which fills 210 ml coffee in each cup, on average. You are supposed to perform find some statistics as a part of quality control department. If the amount of drink dispensed regularly from this machine is normally distributed with a standard deviation equal to 13.5 ml,

 

a) What percentage of the cups will contain more than 231 ml?

b) What would be the chance that a cup will contain between 180 and 220 ml?

c) Calculate the number of cups that will have a probability to overflow if 235 ml cups are used for 900 drinks sold at any random day?

d) If only less than 15% cups are overflowed, what would the cutoff filling volume be?




1
Expert's answer
2021-08-10T12:26:03-0400

Let "X=" the amount of coffee in a cup: "X\sim N(\mu, \sigma^2)."

Given "\mu=210\ ml, \sigma=13.5\ ml."

a)


"P(X>231)=1-P(Z\leq\dfrac{231-210}{13.5})"

"=1-P(Z\leq \dfrac{14}{9})\approx0.0599"

6% of the cups will contain more than 231 ml.


b)


"P(180<X<220)=P(X<220)-P(X\leq180)"

"=P(Z<\dfrac{220-210}{13.5})-P(Z\leq\dfrac{180-210}{13.5})"




"\approx0.7705747-0.0131341\approx0.757441"

"0.757441"


c)

"P(X>235)=1-P(Z\leq\dfrac{235-210}{13.5})"

"=1-P(Z\leq\dfrac{50}{27})\approx0.032024"

"0.032024(900)=29"

"29" cups


d)


"P(X>x)=0.15"


"P(X>x)=1-P(Z\leq\dfrac{x-210}{13.5})=0.15"

"P(Z\leq\dfrac{x-210}{13.5})=0.85"

"\dfrac{x-210}{13.5}\approx1.036433"

"x=224"

"224" ml.